# Quantum Geometry: Revisiting electronic scales in quantum matter

### v1

@@ IV.2 Transport coefficients
Non-reciprocal conductivity $\sigma^{(2)}_{\mu\mu\nu}$ $\ell_{g}^{3}/\omega_{g}$ $\frac{e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum_{n}f_{n}\left(\partial_{\mu}G^{n}_{\mu\nu}-2\partial_{\nu}G^{n}_{\mu\mu}\right)$ [64, 65, 66, 67, 68, 69]
Optical transition rate $\Gamma_{\mu\nu}(\omega)$ $\ell_{g}^{2}/\omega_{\rm ext}$ $\pi\int_{\rm BZ}\sum^{\prime}_{nm}g^{mn}_{\mu\nu}\delta(\omega-\omega_{mn})$ [70, 71]
@@
Current injection (linear pol.) $\sigma^{(2)}_{\mu\mu\lambda}(\omega)$ $\ell_{\rm mr}\ell_{g}^{3}\omega_{\rm ext}$ $\tau\frac{2\pi e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum^{\prime}_{nm}\partial_{\lambda}\omega_{mn}g_{\mu\nu}^{nm}\delta(\omega-\omega_{mn})$ [72, 73, 74, 75]
@@
Current injection (circular pol.) $\sigma^{(2)}_{\mu\nu\lambda}(\omega)$ $\ell_{\rm mr}\ell_{g}^{3}\omega_{\rm ext}$ $\tau\frac{2\pi e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum^{\prime}_{nm}\partial_{\lambda}\omega_{mn}\Omega_{\mu\nu}^{nm}\delta(\omega-\omega_{mn})$ [76, 73, 74, 75]
@@
Shift current $\sigma^{(2)}_{\mu\nu\lambda}(\omega)$ $\ell_{g}^{3}$ $\frac{i\pi e^{3}}{-2\hbar^{2}}\int_{\rm BZ}\sum^{\prime}_{nm}(Q^{nm}_{\mu\nu\lambda}-Q^{nm*}_{\nu\mu\lambda})\delta(\omega-\omega_{mn})$ [77, 78]
Spectral weight $n/m_{\lambda}$ $\ell_{g}^{2}\omega_{g}$ $(1/\hbar)\int_{\rm BZ}\sum^{*}_{nm}\omega_{mn}g^{nm}_{\lambda\lambda}$ [32]
Orbital magnetization $M_{\lambda}$ $\ell_{g}^{2}\omega_{g}$ $(e/c)\int_{\rm BZ}\sum^{*}_{nm}\omega_{mn}\Omega^{nm}_{\mu\nu}\epsilon_{\mu\nu\lambda}$ [79]
Localization tensor $g_{\lambda}$ $\ell_{g}^{2}$ $\int_{\rm BZ}\sum^{*}_{nm}g^{nm}_{\lambda\lambda}$ [57]
Chern number $\mathcal{C}_{\lambda}$ $\ell_{g}^{2}$ $\epsilon_{\mu\nu\lambda}\int_{\rm BZ}\sum^{*}_{nm}\Omega^{nm}_{\mu\nu}$ [80]
@@
Table 1: Examples of geometric and mixed transport coefficients and some optical responses. Definitions: Occupation function $f_{n}$, difference of band energies $\hbar\omega_{nm}=\hbar\omega_{n}-\hbar\omega_{m}$, and the double sum over occupied and empty states is defined as $\sum^{*}_{nm}=\sum_{n,m\neq n}f_{n}(1-f_{m})$, while $\sum^{\prime}_{nm}=\sum_{n,m\neq n}(f_{n}-f_{m})$. To estimate the units of the response, divide by the volume on the unit cell and consider the units of its prefactors. The symbol $\epsilon_{\mu\nu\lambda}$ is the Levi-Civita totally antisymmetric tensor and $\delta_{mn}$ the Kronecker delta. Only the interband, geometric contributions.
Much progress has been made in diagnosing the most well-known part of the QGT - the Berry curvature ($\ell_{g}^{2}$), which can be recast as a single projection operation. As mentioned before, the Berry curvature leads to the intrinsic anomalous Hall effect, which describes the transverse, antisymmetric current component that arises in magnetic materials, which thus carry a nonzero momentum average of the Berry curvature. In magnetic insulators, the same mechanism gives rise to the quantum anomalous Hall effect [58, 59]. More recently, high harmonic generation and polarization-sensitive ARPES techniques were able to directly measure the momentum-resolved Berry curvature in SiO2 [81] and WSe2 [82]. Another route to measure the Berry curvature is to identify observables which couple to its momentum derivatives. For example, the nonlinear intrinsic anomalous Hall effect has been successfully expressed in terms of the quasiparticle lifetime times the momentum derivative of the Berry curvature ($\ell_{\rm mr}\ell_{g}^{3}$[61, 62, 63]. This derivative has become known as the Berry curvature dipole.
On the other hand, no direct measurements of the momentum-resolved quantum metric ($\ell_{g}^{2}$) or its momentum average have been reported so far. There have been suggestions that the metric might couple directly to the dc-conductivity [83, 84, 85] in certain fine-tuned settings, but no linear response observable is known which is generically determined solely by the quantum metric. Even though the quantum metric of the ground state seems rather elusive, the corresponding dipole matrix elements appear almost ubiquitously in response functions in the form of wavefunction overlaps. These matrix elements may have origin in quantum geometry or simply due to onsite atomic transitions, and are generally weighted by velocities or energies. The resulting observable is thus a mixed quantity which depends on quantum geometry and the band energies. To give one example, in a multi-orbital system, the nonlocal conductivity acquires a complicated dependence on several dispersive and geometric features [86]. Many cases of these mixed responses have been reported, but particularly successful have been examples where the interband quantum metric is normalized by the band gaps ($\ell_{g}^{2}/\omega_{g}$). Such a term accounts for the unavoidable mixing of bands in the presence of an adiabatic perturbation. This is highly analogous to the Landau-Zener effect [87] and is thus most appropriately viewed as the excitation of virtual interband transitions [88]. Landau Zener mixing can be identified in the low frequency linear capacitance of insulators, which has been employed to illuminate the role of the QGT for dielectric properties [36]. A similar effect can be observed in the nonlinear dc-conductivity of time-reversal breaking band structures [89, 75], leading to a novel nonlinear Hall effect and non-reciprocal longitudinal conductivity [64, 65]. Both contributions originate from the dipole of the Landau-Zener mixing ($\ell_{g}^{3}/\omega_{g}$[66, 67, 68, 69]. Recently, the third order conductivity has been derived for the study of altermagnets, which yields structurally analogous mixed response functions with Berry curvature and Landau-Zener terms as building blocks, albeit in more complicated combinations ($\ell_{g}^{4}/\omega_{g}$[90, 91].
@@
Finally, one can search for responses which contain the band-resolved QGT in a given response function. Such is often possible in resonant optical responses. Notably, the interband matrix elements from optical transition rates (i. e. Fermi’s golden rule) yield the QGT, but evaluated with respect to two energies which are separated by the photon energy of the incident light (scale $\ell_{g}^{2}/\omega_{\rm ext}$[70, 71]. Furthermore, in materials breaking time-reversal symmetry, a nonreciprocal directional dichroism has been predicted which is proportional to the quantum metric times the quasiparticle velocity $(\ell_{g})$  [92, 93]. Several examples of resonant mixing have also been reported for the nonlinear optical conductivity [73, 74, 75]. Namely, there is a quantized circular photogalvanic effect in two-band systems originating from the Berry charges located at Weyl points ($\ell_{\rm mr}\ell_{g}^{3}\omega_{\rm ext}$[76]. More generally, shift ($\ell_{g}^{3}$) and injection currents ($\ell_{\rm mr}\ell_{g}^{3}\omega_{\rm ext}$) have been connected to matrix elements of the QGT [72] and to geometric objects which go beyond the QGT, which has been termed Riemannian geometry  [94, 77, 95, 96] or multi-state geometry [97, 78].
@@
Despite the tremendous progress in the last few years in characterizing geometric features in observables, it is important to stay parsimonious in applying this label. Matrix overlaps are ubiquitous in multi-orbital systems, and not every occurrence of a matrix element automatically signals that quantum geometry (dipole fluctuations from orbital frustration and delocalization) is an important ingredient towards understanding a particular response. We should be aware of the fact that the momentum structure of the matrix overlaps leads in many cases only to quantitative modifications, without changing the overall character of the response. The examples listed above show that it is nonetheless possible to distill and isolate the role of quantum geometry, and we are confident that many more examples like these will be found. However, to reiterate, wave function overlaps are a necessary, but not a sufficient ingredient for quantum geometric effects.
@@
A second important caveat is the fact that almost all of the response functions discussed above can additionally receive large contributions from extrinsic origin, like impurity scattering, phonons and boundary effects [98, 63, 99]. It is an ongoing challenge to disentangle the latter from the intrinsic phenomena which originate from the bulk wavefunctions, and this is only possible on a case-by-case basis.
IV.3 Correlated phases
@@
The SWM sum rule and the localization tensor have been known for over three decades. More recently, the surge of interest in quantum geometry has been largely driven by the discovery of the magic angle in twisted bilayer graphene (TBG) [100, 101, 102, 103] and the rapid development of moiré materials [104]. These systems exhibit flat bands due to destructive interference of electronic paths over large unit cells, where the geometric length scale is typically comparable to the unit cell size. While moiré materials are expected to host significant quantum geometry, much of the literature on quantum geometric effects differs conceptually from the theme of this review.
Our focus has been on dipole fluctuations in the ground state, whereas in moiré systems, the interest lies in the role of such fluctuations in modifying interactions that drive many-body phases. This arises because quantum geometry derives from dipole fluctuations which smear the real-space charge distribution, causing even short-range interactions to develop spatial profiles that can lead to a completely different ground state. Practically, these effects manifest through form factors, $\langle u_{k}|u_{k+q}\rangle$, in projected interactions, $\sum_{q}V_{q}\bar{\rho}_{q}\bar{\rho}_{-q}$ where the projected density operator is $\bar{\rho}_{q}=\sum_{k}\langle u_{k}|u_{k+q}\rangle c^{\dagger}_{k+q}c^{\phantom{\dagger}}_{k}$. Numerous studies on TBG have already highlighted the importance of form factors within a Hartree-Fock framework. There has also been intense effort to interpret these form factors in terms of well-understood systems like Landau levels [105] and heavy fermions [106].
The finite-$q$ overlaps that appear in projected density are fundamentally different from infinitesimal-$q$ overlaps that appear in projected position operator and the quantum geometric tensor [107]. As a result, while it all too common to find quantum geometric effects in projected interactions from form-factors, it is incredibly difficult to identify an observable that isolates the quantum geometric tensor. On dimensional grounds, such an observable must incorporate an additional length scale to account for the extra factor of $q$ that relates the projected density and position operators, $\bar{\rho}_{q}\approx 1+iq_{\mu}\cdot\bar{r}_{\mu}$.
The $f$-sum rule, defined in Eq. (8) with $\eta=1$, is one such observable with the right dimensions [38]. However, as noted in the previous section, the sum rule carries energetic pre-factors that make it different from the quantum metric. This is where interactions can come into play and replace the pre-factors with an interaction energy scale. While this argument is purely based on dimensional grounds, it is the working principle in flat-band superconductors.
A flat-band superconductor with attractive Hubbard interactions, $V_{q}=-U$, and the uniform pairing condition [108] is described by the Hamiltonian $H=-U\sum_{q}\bar{\rho}_{q}\bar{\rho}_{-q}$ [109]. The low-energy $f$-sum rule is highly sensitive to the quantum geometry of the non-interacting bands [44, 45]. In gapped BCS superconductors at $T=0$, it is directly related to the superfluid stiffness [110], which, within the mean-field approximation [111], is given by $D_{s}\sim n_{\rm 2D}U\ell_{g}^{2}$ where $n_{\rm 2D}$ is the charge density and $\ell_{g}^{2}$ is given by the trace of the quantum metric ${\rm tr}[g]$. Since the superfluid stiffness determines the effective mass of Cooper pairs, the geometric scale renormalizes the Cooper pair mass as $\hbar^{2}/m_{\rm eff}\sim U\ell_{g}^{2}$ [112].
Other than the sum rule and superfluid stiffness, several other observables, including coherence length, have been shown to exhibit quantum geometric corrections in a Landau-Ginzburg theory [113]. Although this framework of mass renormalization has been extensively applied to TBG [114, 115, 116], it is not limited to attractive interactions or superconducting ground states. Similar effects appear in systems with repulsive interactions, where the ground state includes excitons [117, 22, 118, 119].
The insights from TBG are now being extended to transition metal dichalcogenide (TMD) heterostructures, which exhibit a wide range of phenomena, including superconductivity [120], fractional quantum Hall phases [121, 122, 123, 124], quantum anomalous Hall effects [125], generalized Wigner crystals [126, 127], and various other correlated phases [128, 129, 130, 131, 132]. While form factors in projected densities are crucial for accurately modeling interactions [133], a comprehensive discussion of this growing body of work is beyond the scope of this review. For broader coverage of the field, we refer readers to [134].
We note in passing that moiré materials are not the only correlated systems where quantum geometric effects are apparent. Excitons in TMDs have long been observed to deviate from the effective mass approximation [135], with some of these deviations attributed to quantum geometric contributions [136, 137, 138]. Lastly, it is also becoming increasingly evident that a correlated ground state is not necessary for quantum geometric effects. The inter-band component of the Coulomb interaction can itself conspire with the inter-band current operator to renormalize the effective mass of carriers [139, 140]. This renormalization manifests in the plasmon group velocity of metals, which can be measured experimentally [141].
V Discussion and Outlook
The importance of quantum interference and orbital mixing across distinct lattice sites cannot be overstated. These quantum effects introduce electron dynamics at scales comparable to (and sometimes larger than) the lattice constant, fundamentally modifying the physical properties and phases of quantum materials.
@@ V Discussion and Outlook
Quantum geometry provides the mathematical framework to systematically account for these position fluctuation effects, both through matrix elements in correlation functions and through form factors in projected electron-electron interactions. As we emphasize in this overview, these effects can be understood by the length and time scales which are imposed not by the band dispersion, but by the momentum dependence of the electron wavefunctions. Therefore, we advocate to revisit minimal models of quantum materials from this fresh perspective. We should start by estimating the magnitude of geometric effects in materials with exotic physical properties. As described here, these scales can be quickly estimated using the optical conductivity (see Fig.3) both in ab-initio computations and direct measurement, and complemented by various other observables organized in Table 1.
That said, many fundamental questions remain open, and the full implications of quantum geometry in solid state physics are in its infancy. Much of the focus on quantum geometric effects have been given to insulators and flat bands [108], when in fact these are universal effects that dictate the impact of having multiple orbitals. Metals may be the most natural setting where quantum geometric effects can manifest, reformulating the conventional expectations of the response and correlated instabilities of various materials. Let us here list several key questions and directions we hope to see developments in the future:
Better models.
@@
Understanding the quantum geometry of materials and how it emerges by specific lattice motifs can guide the search for materials with enhanced responses or novel quantum phases. Large geometric effects often signal the presence of exotic physics, so how can we utilize the vast computational capabilities of material science to espouse geometric insights? Can we guess which materials will have which order based not only on fermiology but also on geometry?
@@
A key open challenge is understanding how interactions combined with significant dipole fluctuations affect the dynamics of electrons around the Fermi surface in metals. Can these effects be quantified through electronic transport at the Fermi surface, such as in its quantum oscillations? How can interactions with energies substantially smaller than the geometric energy $U\ll\hbar\omega_{g}$ lead to the transfer of spectral weight to zero frequencies, enhancing charge transport, both in metals [141] and correlated condensates [108, 22]? What is the mechanism by which geometry “speeds up” charge carriers in quantum matter?
Interpretation of experiments
Quantum geometry can help interpret experimental data, particularly when different probes appear to measure different effective masses or when responses deviate from semiclassical expectations.
@@
Landau levels have extraordinary properties and are known to stabilize many exotic quantum states such as fractional topological order, exciton condensates [142]. However, it comes at the expense of high magnetic fields. The geometric framework introduces a geometric frequency $\omega_{g}$ and a geometric length $\ell_{g}$ that mimic the cyclotron frequency and magnetic lengths, however these emerge from lattice interference and not from the external Lorentz force. With the recent advent of moiré heterostructures, many quantum Hall phenomena have been mimicked in the absence of a magnetic field, opening a myriad of possibilities for new quantum devices. However, moiré systems come with their own problem of scalability. The proof of concept has been done - can we now find scalable materials with exotic orders which can be incorporated in quantum devices?

### v2

@@ Introduction
Non-reciprocal conductivity $\sigma^{(2)}_{\mu\mu\nu}$ $\ell^{3}/\mathcal{E}$ $\frac{e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum_{n}f_{n}\left(\partial_{\mu}G^{n}_{\mu\nu}-2\partial_{\nu}G^{n}_{\mu\mu}\right)$ [80, 81, 82, 83, 84, 85]
Optical transition rate $\Gamma_{\mu\nu}(\omega)$ $\ell^{2}/\omega$ $\pi\int_{\rm BZ}\sum^{\prime}_{nm}g^{mn}_{\mu\nu}\delta(\omega-\omega_{mn})$ [86, 87]
@@
Current injection (linear pol.) $\sigma^{(2)}_{\mu\mu\lambda}(\omega)$ $\lambda\mathcal{V}\ell^{2}$ $\tau\frac{2\pi e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum^{\prime}_{nm}\partial_{\lambda}\omega_{mn}g_{\mu\nu}^{nm}\delta(\omega-\omega_{mn})$ [88, 89, 90, 91]
@@
Current injection (circular pol.) $\sigma^{(2)}_{\mu\nu\lambda}(\omega)$ $\lambda\mathcal{V}\ell^{2}$ $\tau\frac{2\pi e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum^{\prime}_{nm}\partial_{\lambda}\omega_{mn}\Omega_{\mu\nu}^{nm}\delta(\omega-\omega_{mn})$ [92, 89, 90, 91]
@@
Shift current $\sigma^{(2)}_{\mu\nu\lambda}(\omega)$ $\ell^{3}$ $\frac{i\pi e^{3}}{-2\hbar^{2}}\int_{\rm BZ}\sum^{\prime}_{nm}(Q^{nm}_{\mu\nu\lambda}-Q^{nm*}_{\nu\mu\lambda})\delta(\omega-\omega_{mn})$ [93, 94]
Spectral weight $n/m_{\lambda}$ $\ell^{2}\mathcal{E}$ $(1/\hbar)\int_{\rm BZ}\sum^{*}_{nm}\omega_{mn}g^{nm}_{\lambda\lambda}$ [22]
Orbital magnetization $M_{\lambda}$ $\ell^{2}\mathcal{E}$ $(e/c)\int_{\rm BZ}\sum^{*}_{nm}\omega_{mn}\Omega^{nm}_{\mu\nu}\epsilon_{\mu\nu\lambda}$ [95]
Many-body metric $g_{\lambda}$ $\ell^{2}$ $(\hbar/\pi e^{2})\int_{\rm BZ}\sum^{*}_{nm}g^{nm}_{\lambda\lambda}$ [66]
Chern number $\mathcal{C}_{\lambda}$ $\ell^{2}$ $\epsilon_{\mu\nu\lambda}\int_{\rm BZ}\sum^{*}_{nm}\Omega^{nm}_{\mu\nu}$ [96]
@@
Table 2: Examples of mixed transport coefficients and some optical responses. Definitions: Occupation function $f_{n}$, and the double sum over occupied and empty states is defined as $\sum^{*}_{nm}=\sum_{n,m\neq n}f_{n}(1-f_{m})$, while $\sum^{\prime}_{nm}=\sum_{n,m\neq n}(f_{n}-f_{m})$. The symbol $\epsilon_{\mu\nu\lambda}$ is the Levi-Civita totally antisymmetric tensor and $\delta_{mn}$ the Kronecker delta. The “Ingredients” column now includes velocities $\mathcal{V}$ in addition of the intrinsic geometric length scale $\ell$ and $\mathcal{E}$. We also highlight external factors such as as an external frequency $\omega$ dictating resonant responses or a momentum relaxation mean free path $\lambda$ due to disorder or interactions. Only the interband, geometric contributions.
Even though the quantum metric of the ground state seems rather elusive, the corresponding dipole matrix elements appear almost ubiquitously in response functions in the form of wavefunction overlaps. These matrix elements may have origins in quantum geometry or simply due to onsite atomic transitions, and are generally weighted by velocities or energies. The resulting observables thus depend on both quantum geometry and band energies. In these cases we must introduce a resonance energy scale $\mathcal{E}$ defined by the band dispersion of the material, as it was done in the previous section. To give one example, in a multi-orbital system, the nonlocal conductivity acquires a complicated dependence on several dispersive and geometric features [97].
Many cases of these mixed responses have been reported, but particularly successful have been examples where the interband quantum metric is normalized by the band gaps ($\ell^{2}/\mathcal{E}$). Such a term accounts for the unavoidable mixing of bands in the presence of an adiabatic perturbation. This is highly analogous to the Landau-Zener effect [98] and is thus most appropriately viewed as the excitation of virtual interband transitions [99]. Landau Zener mixing can be identified in the low frequency linear capacitance of insulators, which has been employed to illuminate the role of the QGT for dielectric properties [41].
A similar effect can be observed in the nonlinear dc-conductivity of time-reversal breaking band structures [100, 91], leading to a novel nonlinear Hall effect and non-reciprocal longitudinal conductivity [80, 81]. Both contributions originate from the dipole of the Landau-Zener mixing ($\ell^{3}/\mathcal{E}$[82, 83, 84, 85]. Recently, the third order conductivity has been derived for the study of altermagnets, which yields structurally analogous mixed response functions with Berry curvature and Landau-Zener terms as building blocks, albeit in more complicated combinations ($\ell^{4}/\mathcal{E}$[101, 102].
@@
Finally, one can search for responses which contain the band-resolved QGT in a given response function. This is often possible in resonant optical responses, defined by an external photon frequency $\omega$ of the incident light. Notably, the interband matrix elements from optical transition rates (Fermi’s golden rule) yield the QGT, but evaluated with respect to two energies which are separated by the photon energy of the incident light ($\ell^{2}/\omega$[86, 87]. Several examples of resonant mixing have also been reported for the nonlinear optical conductivity [89, 90, 91]. Namely, there is a quantized circular photogalvanic effect (cf. circular polarized current injection in Table 2) in two-band systems originating from the Berry charges located at Weyl points [92], whose size $\lambda\ell^{2}$ is the result of a convenient cancellation of the Fermi velocities with the integration measure. More generally, shift ($\ell^{3}$) and injection currents ($\lambda\mathcal{V}\ell^{2}$) have been connected to matrix elements of the QGT [88] and to geometric objects which go beyond the QGT, which has been termed Riemannian geometry  [103, 93, 104, 105] or multi-state geometry [106, 94]. Experimentally, bulk photovoltaic responses have been prominently observed in semimetals like CoSi [107] and TaAs [108, 109].
As an illustration, let us analyze the the longitudinal injection current (Table 2). It features three ingredients, (i) the difference of the band velocities between valence and conduction band, (ii) a delta function which enforces resonance and (iii) the matrix elements of the interband quantum metric. It is therefore possible to combine (i) and (ii) in order to write the injection current as the integral over the quantum metric over the space of optically allowed transitions $\bm{S}_{k}$,
$\sigma^{(2)}_{\mu\mu\nu}$ $=-\tau\frac{2\pi e^{3}}{\hbar^{2}}\hat{\nu}\cdot\int_{\omega_{mn}=\omega}\!\!\!d\bm{S}_{k}g^{\mu\mu}.$ (9)
As discussed before, this integral measures properties of the interband quantum geometry, not the ground state metric itself.
The overall scaling of the response, $\sigma^{(2)}_{\mu\mu\nu}\approx\lambda\mathcal{V}\ell^{2}$, can be understood from three physical ingredients. First, the response is proportional to the mean free path $\lambda$ because the injected carriers must scatter to form a steady DC current. Second, the strength of the resonant optical transition, driven by a photon of frequency $\omega$, is governed by the quantum metric, which provides an area scale $\ell^{2}$. Third, a net current requires an asymmetry in the band structure, provided by the velocity difference between bands, which contributes the final intrinsic scale $\mathcal{V}$.
Injection currents require a mismatch in band velocities between time-reversed processes; an intuitive interpretation can therefore be given in terms of the two triangle diagrams that constitute this response, which are related by time reversal [88, 91]. In the absence of time reversal, the semiclassical quasiparticle motion in the clockwise and anti-clockwise nonlinear excitation does not occupy identical real-space areas because they involve different velocities. The effect is thus proportional to the transition area $\ell^{2}$ times the intrinsic velocity mismatch, which depends on the external frequency $\mathcal{V}$.
@@
Despite the tremendous progress in the last few years in characterizing geometric features in observables, it is important to stay parsimonious in applying this label. Matrix overlaps are ubiquitous in multi-orbital systems, and not every occurrence of a matrix element automatically signals that quantum geometry (dipole fluctuations from orbital frustration and delocalization) is an important ingredient towards understanding a particular response. We should be aware of the fact that the momentum structure of the matrix overlaps leads in many cases only to quantitative modifications, without changing the overall character of the response.
@@
Nonetheless, the examples listed above show that it is possible to distill and isolate the role of quantum geometry, and we are confident that many more examples like these will be found. Another important aspect is that almost all of the response functions discussed above can additionally receive large contributions from extrinsic origin ($\lambda$), like impurity scattering, phonons and boundary effects [110, 73, 111]. It is an ongoing challenge to disentangle the latter from the intrinsic phenomena ($\ell$) which originate from the bulk wavefunctions, and this is only possible on a case-by-case basis.
Correlated phases
@@
The SWM sum rule and the localization tensor have been known for over three decades. More recently, the surge of interest in quantum geometry has been driven by the discovery of the magic angle in twisted bilayer graphene [112, 113, 114, 115] and the rapid development of moiré materials [116]. These systems exhibit flat bands due to destructive interference over large moiré unit cells $a_{m}$, for which the geometric length scale is typically $\ell_{g}\approx a_{m}\gg a$. While moiré materials clearly host significant geometric effects, much of the existing literature pursues questions that differ conceptually from the theme of this review. Our focus has been on ground-state dipole fluctuations, whereas in moiré systems the interest lies in how such fluctuations reshape interactions and thereby stabilize correlated phases.
Nevertheless, viewing the problem through the lens of a geometric length scale provides useful intuition. Dipole fluctuations are tied to extended real-space charge distributions, which can cause even short-range interactions to acquire spatial structure at the geometric scale. In practice, these effects enter through projected interactions,
$\bar{H}_{\rm int}=\sum_{q}V_{q}\bar{\rho}_{q}\bar{\rho}_{-q},\qquad\bar{\rho}_{q}=\sum_{k}\langle u_{k}|u_{k+q}\rangle\,c^{\dagger}_{k+q}c_{k},$ (10)
where the form factors $\langle u_{k}|u_{k+q}\rangle$ encode the quantum geometry.
A central conceptual point is that projected interactions depend on the form-factor structure at all $q$. This differs fundamentally from the small-$q$ overlaps that define the projected position operator and the quantum geometric tensor [117]. Because interactions probe higher moments of the dipole operator, isolating contributions that depend only on the small-$q$ geometric tensor is generally difficult. By dimensional analysis, any observable depending solely on the quantum metric must introduce an additional length scale to compensate for the extra factor of $q$ relating projected density to projected position.
A remarkable example where small-$q$ geometry becomes decisive is the stabilization of lattice fractional Chern insulators. The form factors provide a direct route to understanding this mechanism [118]. A key result is the trace condition [28], which fixes the relation between the metric and Berry curvature and constrains fluctuations of the vortex operator $z=x+iy$. Satisfying this constraint is sufficient for obtaining exact fractional Chern insulator ground states for short-range pseudopotentials, in analogy with Laughlin’s construction [119, 120]. Related notions, such as ideal quantum geometry [121], Kähler geometry [122], and vortexability [119], are alternative formulations that have gained prominence because in the chiral limit of twisted bilayer graphene [123], the wavefunctions share similarities with the lowest Landau level [124, 120]. The broader role of geometry in fractional Chern insulators continues to develop, especially following recent observations of fractional states in moiré systems [125, 126, 127].
Another setting where geometric effects manifest cleanly is flat-band superconductivity. Based solely on dimensions, the $f$-sum rule of Eq. (8) (with $\eta=1$) is a good candidate for quantum geometric contributions. Although the sum rule carries energetic prefactors unrelated to the quantum metric, interactions can replace these with an interaction energy scale, enabling a finite low-energy sum rule even in a perfectly flat band. Because the superfluid stiffness is bounded above by this sum rule [128], a nonzero stiffness becomes possible in flat bands.
At a microscopic level, however, the superfluid stiffness generally depends on the full form-factor structure of the projected interaction, not solely on small-$q$ dipole fluctuations. Although this makes the problem sensitive to model-specific details, a particularly clear illustration was provided in Ref.[129], which showed that in flat-band lattice models, even a perfectly flat band can support a finite superfluid stiffness when attractive interactions and quantum geometry are both present. In their calculation for a multi-orbital attractive Hubbard model with the uniform pairing condition [130] and within single-mode approximation, the stiffness scales as $D_{s}\propto nU\ell^{2}$, where $n$ is the charge density and $\ell^{2}$ is a geometric length scale given by the trace of the quantum metric of the flat band [131, 132]. Using projection based on Wannier functions, it was shown that this effect is a consequence of projected interactions containing a pair-hopping process with effective charge $2e$, whose strength is highly sensitive to the underlying quantum geometry [133]. Subsequent works reformulated this phenomenon in terms of position operators, emphasizing that unlike the full density operator, the projected density couples to external gauge fields and hence leads to a non-vanishing low-energy optical response [51, 52, 53].
The precise scaling between the superfluid stiffness and the geometric length scale can be refined using the generalized random phase approximation [134], which preserves the same functional form but includes the minimal quantum metric [135, 136, 137]. Physically, the stiffness sets the effective Cooper-pair mass, so the relation $1/m_{\rm eff}\approx U\ell^{2}$, can be viewed as the geometric length scale renormalizing the pair mass, as confirmed from a calculation of the two-body bound state with separable interaction–potentials [138]. Quantitative estimates of this renormalization have been obtained in twisted bilayer graphene [6, 7, 5]. Beyond stiffness, several other superconducting properties exhibit quantum geometric contributions, including the coherence length [139], the Josephson current [140], and the localization of Majorana zero modes [141].
More broadly, the interplay of quantum geometry and correlations is not restricted to fractional Chern insulators or superconducting ground states. Similar effects appear as well in systems where the ground state supports excitons [142, 3, 143, 144] or magnetic order, with geometry contributing to flat-band ferromagnetism [145] and spin stiffness [146]. Finally, we emphasize that a correlated ground state is not a prerequisite: even in weakly interacting bands the interband Coulomb interaction, together with the interband current operator, can renormalize the carrier effective mass [147, 148, 117] and optical conductivity [149, 150], though this is yet to be observed experimentally.
We note that a large geometric length scale does not always lead to exotic correlated phases or unusual transport. We have in this review focused on $\ell_{g}$ appearing from lattice interference at the unit cell scale, but polarization fluctuations may appear through other sources of interference. Metals with impurities provide a simple illustration[151]. Impurities introduce an effective random lattice that reduces the quantum metric from infinity to a large finite value given by the mean free path, $\lambda\gg a$, yet this change leaves the metallic state essentially unchanged and often well described withing semiclassical approximations and Fermi liquid theory. Higher Landau levels offer another counter-example in a similar vein. The quantum metric grows with the Landau level index $N$, but at large $N$, the system enters a semiclassical regime where bubble and stripe phases appear rather than fractional Chern insulators [152, 153]. Therefore we should be careful to evaluate in every case whether geometry is reshaping phenomena in a meaningful way.
Outlook
@@
Quantum interference and orbital mixing across distinct lattice sites introduce electron dynamics at scales comparable to (and sometimes larger than) the lattice constant, fundamentally modifying the physical properties and phases of quantum materials. Quantum geometry provides a systematic framework for capturing the phenomenology arising from dipole fluctuations. As we discussed above, these effects can be understood by the length and time scales which are imposed not by the band dispersion, but by the momentum dependence of the electron wavefunctions. Therefore, we advocate revisiting minimal models of quantum materials from this fresh perspective. We should start by estimating the magnitude of geometric effects in materials with exotic physical properties. As described here, these scales can be quickly estimated using the optical conductivity (Fig.3) both in ab-initio computations and direct measurement, and complemented by various other observables organized in Table 2.
That said, many fundamental questions remain open, and the full implications of quantum geometry in solid state physics are in its infancy. Much of the focus on quantum geometric effects have been given to insulators and flat bands [132], when in fact these are universal effects that dictate the impact of having multiple orbitals. Metals may offer the next natural setting to explore quantum geometric effects, where they could play a role in driving instabilities toward a particular correlated phase. While the ground-state metric diverges in a metal, recent works particularly in kagome metals [154], suggests that nontrivial wavefunction geometry along the Fermi surface can strongly influence susceptibilities to ordered phases. This motivates a re-examination of how quantum geometry might renormalize electron dynamics at the Fermi surface, although its precise connection to traditional frameworks like Landau’s Fermi liquid theory remains an open question.
Here we list several key questions and directions we hope to see developments in the future:
Better models
@@
Understanding the quantum geometry of materials and how it emerges by specific lattice motifs can guide the search for materials with enhanced responses or novel quantum phases. How can we utilize the vast computational capabilities of material science to espouse such geometric insights? Can we guess which materials will have which order based not only on fermiology but also on geometry?
@@
A key open challenge is understanding how interactions combined with significant dipole fluctuations affect the dynamics of electrons around the Fermi surface in metals. Can these effects be quantified through electronic transport at the Fermi surface, such as in its quantum oscillations? It is now recognized that even weak interactions $U\ll\mathcal{E}$ can transfer spectral weight from high frequencies ($\hbar\omega\gg\mathcal{E}$) to low frequencies, thereby enhancing charge transport, both in metals [155] and in correlated condensates [132, 3]. However, whether this redistribution reflects a universal geometric mechanism remains an open question. In other words, can quantum geometry universally “speed up” charge carriers in quantum matter? If so, what experiments could isolate carrier dynamics with and without geometric contributions?
@@
Landau levels have extraordinary properties and are known to stabilize many exotic quantum states such as fractional topological order or exciton condensates [156]. However, it comes at the expense of high magnetic fields. The geometric framework introduces a natural frequency $\mathcal{E}/\hbar$ and an intrinsic geometric length $\ell$ that mimic the cyclotron frequency and magnetic lengths, however these emerge from lattice interference and not from the external Lorentz force. With the recent advent of moiré heterostructures, many quantum Hall phenomena have been mimicked in the absence of a magnetic field, opening a myriad of possibilities for new quantum devices. However, moiré systems come with their own problem of scalability. The proof of concept has been done - can we now find scalable materials with exotic orders which can be incorporated in quantum devices?